# Topology of dyonic AdS black holes with quasitopological electromagnetism in Einstein-Gauss-Bonnet gravity

### v2

@@ Abstract
Abstract: In this study, we employ the thermodynamic topological method to classify critical points for the dyonic AdS black holes with quasitopological electromagnetism in the Einstein-Gauss-Bonnet background. To this end, we find a small/large black hole phase transition in all dimensions of space-time, the existence of a conventional critical point implies a total topological charge of $Q_{t}=-1$. The coupling constant $\alpha$ gives rise to a more intricate phase structure, with the emergence of a triple points requires $\alpha\geq 0.5$ and $d=6$. Interestingly, the condition for the occurrence of small/intermediate/large phase transition is that the coupling constant a takes a special value ($\alpha=0.5$), the two conventional critical points $CP_{1},CP_{2}$ of the black hole are physical critical point, and the novel critical point $CP_{3}$ that lacks the capability to minimize the Gibbs free energy. The critical points $CP_{1}$ and $CP_{2}$ are observed to occur at the maximum extreme points of temperature in the isobaric curve, while the critical point $CP_{3}$, emerges at the minimum extreme points of temperature. Furthermore, the number of phases at the novel critical point exhibits an upward trend, followed by a subsequent decline at the conventional critical points. With the increase of the coupling constant ($\alpha=1$), although the system has three critical points, only the conventional $CP_{1}$ is a (physical) critical point, and the conventional $CP_{2}$ serves as the phase annihilation point. This means that the coupling constant $\alpha$ has a non-negligible effect on the phase structure.
@@ I Introduction
The official announcement on February 11, 2016 of the detection of gravitational waves by the LIGO represents a groundbreaking achievement that provides significant validation for Einstein’s general relativity. This momentous discovery not only confirms the existence of black holes but also verifies the presence of gravitational waves ch1 ; ch2 ; ch3 ; ch4 . The emergence of black hole (BH) thermodynamics stems from the application of general laws in thermodynamics to the general relativity, with the objective of investigating the characteristics inherent to black holes. The fundamental concept involves regarding black holes as a system governed by thermodynamic principles, wherein mass corresponds to energy, entropy is directly proportional to the surface area, and temperature is represented by surface gravity. In 1974, Hawking made a groundbreaking discovery that black holes exhibit thermal radiation ch5 , this revelation not only resolves the inherent contradictions within the thermodynamics of black holes but also successfully establishes a profound connection between gravitational theory, quantum mechanics, and thermodynamics. The phenomenon of phase transition is commonly observed in thermodynamic systems, wherein the Schwarzschild black hole, when considered as a thermodynamic system within the framework of AdS, exhibits an unstable state from a thermodynamic perspective at low temperatures. However, with increasing temperature, the free energy is reduced, and finally the black hole reaches stability. Essentially, the stability of a black hole hinges on the critical temperature known as the Hawking-Page (HP) phase transition ch6 . Based on the AdS/CFT correspondence theory, there has been considerable focus on studying the HP phase transition, which is recognized as a phase transition between confinement and deconfinement in the gauge/gravity duality ch7 .
@@
Kastor and Ray proposed the concept of regarding the cosmological constant $\Lambda$ as a variable thermodynamic pressure, and suggested that the volume of the black hole in AdS space can be considered as its conjugate quantity, while the black hole’s mass is interpreted as the spacetime’s enthalpy, rather than its internal energy ch8 . This notion was further expanded to Lovelock theory and applied to black holes, which demonstrates that Lovelock coupling serves as an extension of the first law of thermodynamics ch9 . Additionally, both the free energy and of Arnowitt-Deser- Misner (ADM) mass AdS black holes were derived ch10 . Furthermore, investigations on dilaton-AdS black holes have revealed that their mass represents the system’s enthalpy ch11 , the phenomenon has led to a surge in research on thermodynamic phase transitions in AdS black holes. For instance, the charged RN-AdS black hole in four-dimensional spacetime was examined by Kubiznak and Mann, who identified a significant transition between small and large black holes, which bears resemblance to the gas-liquid phase transition observed in van der Waals fluids ch17 . The phase transition in nonlinear electromagnetic field was investigated by Hendi and Vahidinia through considering the Einstein gravitational solution under the background of generalized Maxwell theory ch18 . Critical phenomena and critical exponents near the critical point of Gauss-Bonnet black hole were studied by Wei et al, who also examined the phenomenon of phase transition in spin Kerr black holes with a high number of dimensions ch19 ; ch20 . Zou et al. explored P-V critical behavior of Born-Infeld Anti-deSitter black hole and Gauss-Bonnet AdS black hole ch21 ; ch22 . It is well known that the typical first-order phase transition has only one critical point. However, the existence of multiple critical points is observed in certain categories of black holes, such as, the occurrence of reentrant phase transition ch24 ; ch25 , the presence of triple points in dynamic AdS black holes under quasitopological electromagnetism (QTE) ch26 ; ch27 , and the $\lambda$ phase transition observed in Horava gravity ch28 , among others. A complementary aspect to the investigation of black hole criticality is the concept of thermodynamic topology. The Duan ’s $\varphi$-mapping theory proposed is utilized to explore the criticality of black holes, resulting in distinct characteristics of topological charge exhibited by critical points within the thermodynamic space, these critical points can be classified into conventional and novel types cq4 . Furthermore, Wei et al. employed the thermodynamic topological method to investigate thermodynamic properties of black holes, this approach enables analysis of stability by considering the winding number associated with defect we1 . The thermodynamic topology classification of different black holes are found in detail in we2 ; we2.1 ; we3 ; we4 ; we5 ; we7 ; we9 ; we8 ; we10 ; we12 ; we13 ; we14 ; we15 ; we16 ; we17 ; we17.1 ; we17.2 ; we18 .
@@
The theory of extended gravity, which encompasses Einstein-Gauss-Bonnet (EGB) gravity, provides novel perspectives on the characteristics of black holes ch29 ; ch30 ; ch31 ; ch32 ; ch33 ; ch34 ; ch35 . Through an examination of the thermodynamic characteristics of the Born-Infeld AdS black hole within the framework of 4D EGB gravity, it has been observed that there exist three critical points within the black hole, these critical points exhibit numerous intriguing properties when subjected to EGB gravity ch25 . Recently, the QTE is characterized by the square norm of the topological 4-form $F\wedge F$, indicating that the system’s energy-momentum tensor corresponds to a perfect fluid with isotropic properties ch36 . Based on this, the thermodynamic phase transition of dyonic AdS black holes with QTE is discussed in detail. By changing the coupling parameters, a triple points phase structure is observed ch27 . Additionally, through an examination of the interplay between QTE and EGB gravity, Sekhmani et al. derived a solution for dyonic AdS black holes with a high number of dimensions. Furthermore, they conducted a comprehensive analysis on how the coupling constant $\alpha$ influences both the energy emission rates and black hole shadow ch37 . On this basis, a comprehensive analysis is conducted on the phase transition of the dyonic AdS black holes with QTE in the EGB background, and it is found that the system parameters will appear triple points in special regions cq3 . In this work, we will use the thermodynamic topological method to investigate the critical points, and further understand the influence of GB coupling constant on the phase structure of dyonic AdS black holes. The structure of the manuscript is as follows: In section II, we give a concise overview of the thermodynamic properties of dyonic AdS black holes with QTE in the EGB background and present a comprehensive explanation of Duan’s $\varphi$-mapping theory. In section III, we will use the thermodynamic topology method to classify the critical points in different space-time dimensions. In section IV, we determine the characteristics of the critical point by analyzing the isobaric curve. Section V contains the remarks and conclusions.
@@ II Dyonic AdS black holes and topology
We first consider the d-dimensional action involving EGB gravity and QTE which can be formulated in a minimal coupling manner ch37
@@
wherein, $\Lambda=-\frac{(d-1)(d-2)}{2l^{2}}$ represents the cosmological constant within the AdS space, $\alpha$ is the symbol used to denote the GB coupling constant, which possesses dimensions of [length]${}^{2}$. Here, the GB term can be expressed as
@@
where $p=d-2$, $\mathcal{F}^{2}=\mathcal{F}_{\mu\nu}\mathcal{F}^{\mu\nu}$ and $\mathcal{H}^{2}=\mathcal{H}_{\rho_{1}\ldots\rho_{p}}\mathcal{H}^{\rho_{1}\ldots\rho_{p}}$. Here, $\beta$ means the coupling constant associated with the dimension [length]${}^{2}$. Additionally, the interaction term reads
$\mathcal{L}_{\rm{int}}=\delta_{\gamma_{1}\ldots\gamma_{D}}^{\lambda_{1}\ldots\lambda_{D}}\mathcal{F}_{\lambda_{1}\lambda_{2}}\mathcal{H}_{\lambda_{3}\ldots\lambda_{D}}\mathcal{F}^{\gamma_{1}\gamma_{2}}\mathcal{H}^{\gamma_{3}\ldots\gamma_{D}}.$ (4)
@@
the hypergeometric function, denoted by $\xi$, can be expressed as
@@
$\widetilde{\alpha}=(d-4)(d-3)\alpha.$ (9)
Here, $Q_{e}$ and $Q_{m}$ are essential parameters that indicate the electrical and magnetic charges of black holes, respectively. These parameters are critical in comprehending the characteristics of black holes. $m$ denotes the mass of the solution within a designated parameter domain, serving as the constant of integration derived from the boundary conditions ch37 . In this case, ADM mass is given by
@@
$\mathcal{V}_{d-2}=\frac{2\pi^{\frac{d-1}{2}}}{\Gamma\left(\frac{d-1}{2}\right)},$ (11)
the symbol $\mathcal{V}_{d-2}$ represents the volume of a unit sphere in $(d-2)$ dimensions. The concept of thermodynamic pressure is closely related to the cosmological constant $\Lambda$, specifically, $P=-\frac{\Lambda}{8\pi}$ ch8 . Based on this, the event horizon’s radius can be utilized to ascertain thermodynamic properties of black holes, including Hawking temperature, entropy, volume, and Gibbs free energy
$T$ $=\frac{1}{\left[8(d-2)\pi r_{+}+16\pi r_{+}^{-1}\mathcal{N}_{1}\right]\mathcal{N}_{2}}[(12+12d(d-5)$ (12)
$\left.\left.+32\pi Pr_{+}^{2}+2d(d-5)r_{+}^{-2}\mathcal{N}_{1}-r_{+}^{6-2d}Q_{m}^{2}\right)\mathcal{N}_{2}-r_{+}^{6}Q_{e}^{2}\right],$
with
$\mathcal{N}_{1}$ $=$ $(d-4)(d-3)(d-2)\alpha,$ (13)
$\mathcal{N}_{2}$ $=$ $r_{+}^{2d}+8r_{+}^{4}\beta Q_{m}^{2}\Gamma(d-1)^{2},$ (14)
[…]
$=\frac{\pi^{\frac{d-1}{2}}r_{+}^{d-2}\left[1+2(d-3)(d-2)\alpha r_{+}^{-2}\right]}{2\Gamma\left(\frac{d-1}{2}\right)},$
$V=\left(\frac{\partial H}{\partial P}\right)_{S,Q_{e},Q_{m}}=\frac{2\pi^{\frac{d-1}{2}}r_{+}^{d-1}}{(d-1)\Gamma\left(\frac{d-1}{2}\right)},$ (16)
$G$ $=2(d-2)r_{+}^{2d+2}\mathcal{N}_{3}+2r_{+}^{2d}\mathcal{N}_{1}\mathcal{N}_{3}+\frac{32\pi P\mathcal{N}_{3}r_{+}^{2d+4}}{d-1}$ (17)
$-\frac{\mathcal{N}_{3}}{\left((d-2)r_{+}^{2}+2\mathcal{N}_{1}\right)\mathcal{N}_{2}}\left[2(d-3)(d-2)\alpha+r_{+}^{2}\right]$
×$\times\left[\left(2\mathcal{N}_{4}r_{+}^{2d}-r_{+}^{8}Q_{m}^{2}\right)\mathcal{N}_{2}-r_{+}^{2d+8}Q_{e}^{2}\right]$
$+\frac{r_{+}^{8}\left(Q_{m}^{2}+Q_{e}^{2}\xi_{r_{+}}\right)\mathcal{N}_{3}}{d-3},$
with
$\mathcal{N}_{3}=\frac{\pi^{\frac{d-3}{2}}r_{+}^{-5-d}}{16\Gamma\left(\frac{d-1}{2}\right)},$ (18)
$\mathcal{N}_{4}=6r_{+}^{2}+d(d-5)r_{+}^{2}+16P\pi r_{+}^{4}+(d-5)\mathcal{N}_{1}.$ (19)
@@
$\Psi=\frac{1}{\sin\theta}T\left(S,x^{i}\right),$ (21)
@@
$\phi^{S}=(\partial_{S}\Psi)_{\theta,x_{i}},\phi^{\theta}=(\partial_{\theta}\Psi)_{S,x_{i}}.$ (22)
@@
the winding number $w_{i}$ determines the topological charge $Q$ at zero points $(\phi^{a}(x^{i})=0)$, with negative or positive values corresponding to a winding number of $-1$ or $+1$ respectively. The summation of charges at all critical points within a thermodynamic system defines the total topological charge $Q_{t}$, this fundamental concept holds great significance in the analysis of thermodynamic systems as it aids in comprehending their global characteristics and facilitating classification. By aggregating the charge associated with each critical point, we can derive a quantitative indicator to depict the system’s topology structure. The classification of distinct thermodynamic systems is determined based on the total topological charge. The aforementioned approach will be employed to investigate dyonic AdS black holes with QTE in the context of EGB gravity.
Figure 1: $d=4$: Up panel: The red arrows represent the vector field $n$ for dyonic AdS black holes with QTE in the EGB background. The green dot are $(r,\theta)=(2.449490,\pi/2)$, which represent the critical points $CP_{1}$. Down panel: The deflection angle $\Omega(\vartheta)$ as a function of $\vartheta$ for contours $C_{1}$ (green curve), $C_{2}$ (brown curve).

### v3

# Thermodynamic topology of dyonic AdS black holes with quasitopological electromagnetism in Einstein-Gauss-Bonnet gravity

@@ Abstract
Abstract: In this study, we investigate the thermodynamic topology of the high-dimensional dyonic AdS black holes with quasitopological electromagnetism in the Einstein-Gauss-Bonnet background. We first examine the topological charge connected to the critical point and find that the two conventional critical points $CP_{1},CP_{2}$ of the black hole are physical critical point, and the novel critical point $CP_{3}$ that lacks the capability to minimize the Gibbs free energy ($\alpha=0.5$). The critical points $CP_{1}$ and $CP_{2}$ are observed to occur at the maximum extreme points of temperature in the isobaric curve, while the critical point $CP_{3}$, emerges at the minimum extreme points of temperature. Furthermore, the number of phases at the novel critical point exhibits an upward trend, followed by a subsequent decline at the conventional critical points. With the increase of the coupling constant ($\alpha=1$), although the system has three critical points, only the conventional $CP_{1}$ is a (physical) critical point, and the conventional $CP_{2}$ serves as the phase annihilation point. This means that the coupling constant $\alpha$ has significant impact on the phase structure. Additionally, we regard dyonic AdS black holes as a topological defect within the thermodynamic space, our findings indicate that alterations in pressure can result in the system exhibiting distinct points of generation and annihilation. However, the total topological number of black holes in different dimensions is $1$, the system shares a similar topological classification as the charged RN-AdS black holes. The discovery we have made provides a crucial component in understanding the thermodynamic topology of dyonic AdS black holes.
@@ I Introduction
The official announcement on February 11, 2016 of the detection of gravitational waves by the LIGO represents a groundbreaking achievement that provides significant validation for Einstein’s general relativity. This momentous discovery not only confirms the existence of black holes but also verifies the presence of gravitational waves ch1 ; ch2 ; ch3 ; ch4 . The emergence of black hole (BH) thermodynamics stems from the application of general laws in thermodynamics to the general relativity, with the objective of investigating the characteristics inherent to black holes. The fundamental concept involves regarding black holes as a system governed by thermodynamic principles, wherein mass corresponds to energy, entropy is directly proportional to the surface area, and temperature is represented by surface gravity. In 1974, Hawking made a groundbreaking discovery that black holes exhibit thermal radiation ch5 , this revelation not only resolves the inherent contradictions within the thermodynamics of black holes but also successfully establishes a profound connection between gravitational theory, quantum mechanics, and thermodynamics. The phenomenon of phase transition is commonly observed in thermodynamic systems, wherein the Schwarzschild black hole, when considered as a thermodynamic system within the framework of AdS, exhibits an unstable state from a thermodynamic perspective at low temperatures. However, with increasing temperature, the free energy is reduced, and finally the black hole reaches stability.
@@
Kastor and Ray proposed the concept of regarding the cosmological constant $\Lambda$ as a variable thermodynamic pressure, and suggested that the volume of the black hole in AdS space can be considered as its conjugate quantity, while the black hole’s mass is interpreted as the spacetime’s enthalpy, rather than its internal energy ch8 . This notion was further expanded to Lovelock theory and applied to black holes, which demonstrates that Lovelock coupling serves as an extension of the first law of thermodynamics ch9 . Additionally, both the free energy and of Arnowitt-Deser-Misner (ADM) mass AdS black holes were derived ch10 . Furthermore, investigations on dilaton-AdS black holes have revealed that their mass represents the system’s enthalpy ch11 , the phenomenon has led to a surge in research on thermodynamic phase transitions in AdS black holes. For instance, the charged RN-AdS black hole in four-dimensional spacetime was examined by Kubiznak and Mann, who identified a significant transition between small and large black holes, which bears resemblance to the gas-liquid phase transition observed in van der Waals fluids ch17 . The phase transition in nonlinear electromagnetic field was investigated by Hendi and Vahidinia through considering the Einstein gravitational solution under the background of generalized Maxwell theory ch18 . Critical phenomena and critical exponents near the critical point of Gauss-Bonnet black hole were studied by Wei et al, who also examined the phenomenon of phase transition in spin Kerr black holes with a high number of dimensions ch19 ; ch20 . Zou et al. explored P-V critical behavior of Born-Infeld Anti-deSitter black hole and Gauss-Bonnet AdS black hole ch21 ; ch22 . It is well known that the typical first-order phase transition has only one critical point. However, the existence of multiple critical points is observed in certain categories of black holes, such as, the occurrence of reentrant phase transition ch24 ; ch25 . For the intriguing solution of black hole with quasi-topological electromagnetism (QTE) sci1 , this theory presents a novel higher-order expansion formulated using the bilinear norm of Maxwell’s theory. In certain specific circumstances, the presence of extra terms does not have any impact on the energy-momentum tensor and the Maxwell equation. The presence of triple points in dynamic AdS black holes with QTE ch26 ; ch27 , and the $\lambda$ phase transition observed in Horava gravity ch28 , among others. A complementary aspect to the investigation of black hole criticality is the concept of thermodynamic topology. The Duan’s $\varphi$-mapping theory proposed is utilized to explore the criticality of black holes, resulting in distinct characteristics of topological charge exhibited by critical points within the thermodynamic space, these critical points can be classified into conventional and novel types cq4 . Furthermore, Wei et al. employed the topological method to investigate thermodynamic properties of black holes, this approach enables analysis of stability by considering the winding number associated with defect we1 . Recently, Fang et al. proposed that the winding numbers can be determined by calculating the residue of the isolated first-order pole of the characterized functions constructed using off-shell free energy. This approach effectively enables the classification of black hole topology l11 .
@@
The theory of extended gravity encompasses Einstein-Gauss-Bonnet (EGB) gravity, which was first proposed by Cisterna et al rna1 , by using this model for other constructions, specifically, extended black strings and p-branes and diverse compactifications with Lovelock fluxes rna2 ; rna3 . EGB gravity provides novel perspectives on the characteristics of black holes ch29 ; ch30 ; ch31 ; ch32 ; ch33 ; ch34 ; ch35 . Through an examination of the thermodynamic characteristics of the Born-Infeld AdS black hole within the framework of 4D EGB gravity, it has been observed that there exist three critical points within the black hole, these critical points exhibit numerous intriguing properties when subjected to EGB gravity ch25 . Recently, the QTE is characterized by the square norm of the topological 4-form $F\wedge F$, indicating that the system’s energy-momentum tensor corresponds to a perfect fluid with isotropic properties sci1 . Based on this, the thermodynamic phase transition of dyonic AdS black holes with QTE is discussed in detail. By changing the coupling parameters, a triple points phase structure is observed ch27 . Additionally, through an examination of the interplay between QTE and EGB gravity, Sekhmani et al. derived a solution for dyonic AdS black holes with a high number of dimensions. Furthermore, they conducted a comprehensive analysis on how the coupling constant $\alpha$ influences both the energy emission rates and black hole shadow ch37 . On this basis, a comprehensive analysis is conducted on the phase transition of the dyonic AdS black holes with QTE in the EGB background, and it is found that the system parameters will appear triple points in special regions cq3 . In this work, we will use the topological method to investigate the critical points, and further understand the influence of GB coupling constant on the phase structure of dyonic AdS black holes. The structure of the manuscript is as follows: In section II, we give a concise overview of the thermodynamic properties of dyonic AdS black holes with QTE in the EGB background and present a comprehensive explanation of Duan’s $\varphi$-mapping theory and the off-shell method. In section III, we will use the topological method to classify the critical points. In section IV, we determine the characteristics of the critical point by analyzing the isobaric curve. In section V.2, we examine dyonic AdS black holes as a defect in the thermodynamic space, aiming to comprehend the local and global topological structure of this defect. Section VI contains the remarks and conclusions.
@@ II Dyonic AdS black holes and topology
We first consider the d-dimensional action involving EGB gravity and QTE which can be formulated in a minimal coupling manner rna1 ; ch37
@@
wherein, $\Lambda=-\frac{(d-1)(d-2)}{2l^{2}}$ represents the cosmological constant within the AdS space, $\alpha$ is the symbol used to denote the GB coupling constant, which possesses dimensions of [length]2. Here, the GB term can be expressed as
@@
where $p=d-2$, $\mathcal{F}^{2}=\mathcal{F}_{\mu\nu}\mathcal{F}^{\mu\nu}$ and $\mathcal{H}^{2}=\mathcal{H}_{\rho_{1}\ldots\rho_{p}}\mathcal{H}^{\rho_{1}\ldots\rho_{p}}$ , in which, according to Ref. rna1 , the gauges fields $\mathcal{F}_{\mu\nu}$ and $\mathcal{H}_{\rho_{1}...\rho_{p}}$ are considered purely electric and magnetic, respectively. Here, $\beta$ means the coupling constant associated with the dimension [length]2. In this case the only non-vanishing terms are $\mathcal{F}_{\mu\nu}\mathcal{F}^{\mu\nu}$, $|\mathcal{H}_{\left[p\right]}|^{2}$, and $|\mathcal{F}\mathcal{H}_{\left[D\right]}|^{2}$. Additionally, the interaction term reads
$\mathcal{L}_{\rm{int}}=\delta_{\gamma_{1}\ldots\gamma_{D}}^{\lambda_{1}\ldots\lambda_{D}}\mathcal{F}_{\lambda_{1}\lambda_{2}}\mathcal{H}_{\lambda_{3}\ldots\lambda_{D}}\mathcal{F}^{\gamma_{1}\gamma_{2}}\mathcal{H}^{\gamma_{3}\ldots\gamma_{D}},$ (4)
@@
here $\alpha$ is the GB coupling constant, the hypergeometric function, denoted by $\xi$, can be expressed as
@@
$\widetilde{\alpha}=(d-4)(d-3)\alpha,$ (9)
wherein, $Q_{e}$ and $Q_{m}$ are essential parameters that indicate the electrical and magnetic charges of black holes, respectively. These parameters are critical in comprehending the characteristics of black holes. $m$ denotes the mass of the solution within a designated parameter domain, serving as the constant of integration derived from the boundary conditions aj1 . In this case, ADM mass is given by
@@
$\mathcal{V}_{d-2}=\frac{2\pi^{\frac{d-1}{2}}}{\Gamma\left(\frac{d-1}{2}\right)}.$ (11)
The equation $f\left(r_{+}\right)=0$ determines the value of $\mathrm{BH}$ horizon radius $r_{+}$, which can be obtained through solving this equation. The concept of thermodynamic pressure is closely related to the cosmological constant $\Lambda$, specifically, $P=-\frac{\Lambda}{8\pi}$ ch8 . Hence, the mass of $\mathrm{BH}$ can be represented in relation to the horizon radius $r_{+}$ as
$M=\frac{1}{32\left(3-4d+d^{2}\right)\pi}r_{+}^{-5-d}\mathcal{V}_{d-2}$ (12)
×$\quad\times\left[(d-1)Q_{m}^{2}r_{+}^{8}+2(d-3)r_{+}^{2d}\right.$
×$\quad\times\left(\left(2-3d+d^{2}\right)r_{+}^{2}\right.$
$\left.\quad+\left(24-50d+35d^{2}-10d^{3}+d^{4}\right)\alpha+16\pi r_{+}^{4}P\right)$
$\quad+(d-1)Q_{e}^{2}r_{+}^{8}{}_{2}F_{1}$
×$\left.\quad\times\left(1,\frac{d-3}{2(d-2)},\frac{7-3d}{4-24},-8\beta Q_{m}^{2}r_{+}^{4-2d}\Gamma(d-1)^{2}\right)\right].$
Furthermore, by utilizing the definition of Hawking temperature as $T=\frac{f^{\prime}\left(r_{+}\right)}{4\pi}$, we can obtain
$T_{+}=\frac{t_{1}}{t_{2}},$ (13)
where
$t_{1}=$ $r_{+}^{-2d-1}\left(8\beta Q_{m}^{2}r_{+}^{4}\Gamma(d-1)^{2}\left(-2r_{+}^{2d}\left(2\Lambda r_{+}^{4}\right.\right.\right.$ (14)
$\left.\left.-(d-3)(d-2)\left(\alpha(d-5)(d-4)+r_{+}^{2}\right)\right)-q_{m}^{2}r_{+}^{8}\right)$
$-r_{+}^{2d+8}\left(Q_{e}^{2}+Q_{m}^{2}\right)-2r_{+}^{4d}\left(2\Lambda r_{+}^{4}\right.$
$\left.\left.-(d-3)(d-2)\left(\alpha(d-5)(d-4)+r_{+}^{2}\right)\right)\right),$
$t_{2}=$ $8\pi(d-2)\left(2\alpha(d-4)(d-3)+r_{+}^{2}\right)$ (15)
×$\times\left(8\beta Q_{m}^{2}r_{+}^{4}\Gamma(d-1)^{2}+r_{+}^{2d}\right).$
The Bekenstein-Hawking entropy in cq3 can be expressed as
$S=\frac{\pi^{\frac{d-1}{2}}r_{+}^{d-2}\left[1+2(d-3)(d-2)\alpha r_{+}^{-2}\right]}{2\Gamma\left(\frac{d-1}{2}\right)},$ (16)
other thermodynamic quantities, including the volume $V$, electric potential $\Phi_{e}$, and magnetic potential $\Phi_{m}$, can be performed by
$V=\frac{2\pi^{\frac{d-1}{2}}r_{+}^{d-1}}{(d-1)\Gamma\left(\frac{d-1}{2}\right)},$ (17)
$\Phi_{e}=\left(\frac{\partial M}{\partial Q_{e}}\right)=\frac{\pi^{\frac{d-3}{2}}r_{+}^{3-d}Q_{e}\xi}{8(d-3)\Gamma\left(\frac{d-1}{2}\right)},$ (18)
$\Phi_{m}$ $=\left(\frac{\partial M}{\partial Q_{m}}\right)$ (19)
$=\frac{\pi^{\frac{d-3}{2}}r_{+}^{3-d}\left[2(d-2)Q_{m}^{2}+(d-3)Q_{e}^{2}\left(r_{+}^{2d}\mathcal{\varsigma}^{-1}-\xi\right)\right]}{16(d-3)(d-2)Q_{m}\Gamma\left(\frac{d-1}{2}\right)},$
where
$\mathcal{\varsigma}=r_{+}^{2d}+8r_{+}^{4}\beta Q_{m}^{2}\Gamma(d-1)^{2}.$ (20)
Similarly, the conjugate quantity to the coupling constant $\alpha$ and $\beta$ can be formulated as
$\Phi_{\alpha}$ $\equiv\left(\frac{\partial M}{\partial\alpha}\right)$ (21)
$=\frac{\pi^{\frac{d-3}{2}}r_{+}^{d-5}\mathcal{\varsigma}_{1}}{8\alpha\Gamma\left(\frac{d-1}{2}\right)}-\frac{(d-3)(d-2)\pi^{\frac{d-1}{2}}r_{+}^{d-4}T}{\Gamma\left(\frac{d-1}{2}\right)},$
where
$\mathcal{\varsigma}_{1}=(d-4)(d-3)(d-2)\alpha$ (22)
$\Phi_{\beta}=\left(\frac{\partial M}{\partial\beta}\right)=\frac{\pi^{\frac{d-3}{2}}r_{+}^{3-d}Q_{e}^{2}\left(r_{+}^{2d}\varsigma-\xi\right)}{32(d-2)\beta\Gamma\left(\frac{d-1}{2}\right)}.$ (23)
By utilizing this information, the generalized Smarr relation is given by
$M$ $=\frac{d-2}{d-3}TS-\frac{2}{d-3}PV+\Phi_{e}Q_{e}+\Phi_{m}Q_{m}$ (24)
$+\frac{2}{d-3}\left(\alpha\Phi_{\alpha}+\beta\Phi_{\beta}\right).$
@@
$\Phi=\frac{1}{\sin\theta}T\left(S,x^{i}\right),$ (26)
@@
$\phi^{S}=(\partial_{S}\Phi)_{\theta,x_{i}},\phi^{\theta}=(\partial_{\theta}\Phi)_{S,x_{i}}.$ (27)
@@
the winding number $w_{i}$ determines the topological charge $Q$ at zero points $(\phi^{a}(x^{i})=0)$, with negative or positive values corresponding to a winding number of $-1$ or $+1$ respectively. The summation of charges at all critical points within a thermodynamic system defines the total topological charge $Q_{t}$, this fundamental concept holds great significance in the analysis of thermodynamic systems as it aids in comprehending their global characteristics and facilitating classification. The exploration of the topology of black holes has been broadened to include various classifications of black holes o1 ; o2 ; o3 ; o4 ; o5 ; o6 ; o7 ; o8 ; o9 ; o10 ; o12 .
A recent study proposes an alternative approach to incorporate topology into the black hole thermodynamics we1 , this perspective implies that defects can be interpreted as black hole solutions. The total winding number, referred to as the topological number, is employed in the classification of distinct black hole solutions. To begin with, we introduce a generalized free energy denoted as $F$, which is represented by
$\mathcal{F}=E-\frac{S}{\tau}.$ (30)
In this context, $E$ and $S$ represent energy and entropy respectively, while $\tau$ is a dimensionless quantity indicating time. The vector field $\phi$ is obtained from
$\phi=\left(\frac{\partial\mathcal{F}}{\partial r_{+}},-\cot\Theta\csc\Theta\right).$ (31)
The vector $\phi$ is centered at $\Theta=\pi/2$, and its unit vector properties are defined by
$n^{a}=\frac{\phi^{a}}{\|\phi\|}\quad(a=1,2)\quad\text{ and }\quad\phi^{1}=\phi^{r_{+}},\quad\phi^{2}=\phi^{\Theta}.$ (32)
The identification of the zero points of $n^{1}$ relies on the specified value of $\tau$, and the calculation of winding numbers is conducted for zero point, the total winding number can be obtained by adding up the individual winding numbers of each branch of the black hole. Up to now, the investigation of topological defects has been expanded to encompass diverse types of black holes sa1 ; ad1 ; ad2 ; l1 ; l2 ; l3 ; l4 ; l5 ; l7 ; l9 ; l10 ; l12 ; l13 ; l14 ; l15 ; l16 ; l16.1 ; l17 ; l18 ; l19 ; l20 ; l21 ; l22 ; l23 ; l24 ; meng1 ; meng2 .